OpenStudy (anonymous):

A firecracker shoots up from a hill 145 feet high with an initial speed of 80 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? a. 13 seconds b. 11 seconds c. 9 seconds d. 6 seconds

OpenStudy (chaise):

I'm pretty sure you're able to sub in the values: s=145 (the y intercept of the graph) and v= 80 (this is how much is changes per a unit) You get: H(t) = -16t^2 + 80t + 145 now you just evaluate the function at t=13, 11, 9, 6 H(t) = -16t^2 + 80t + 145 H(13) = -16*13^2 + 80*13 + 145= H(11) = -16*11^2 + 80*11 + 145= H(9) = -16*9^2 + 80*9 + 145= H(6) = -16*6^2 + 80*6 + 145=

OpenStudy (anonymous):

Thank you

OpenStudy (chaise):

I think the only time the fire cracker is in the air is when t=6. The rest it has well and truly hit the ground, lol. although I could be wrong

OpenStudy (anonymous):

That is what I came up with

OpenStudy (dumbcow):

you could also set H(t) =0 and solve for t using quadratic formula --> x =6.413 just in case its not multiple choice :)